theyadoreshayyy:

What is the length of the support AB? 7 divided by cos 45 degrees 7 tan 45° 7 divided by sin 45 degrees 7 cos 45°

theyadoreshayyy:

new2luv:

can you tell me what you think

theyadoreshayyy:

a

new2luv:

yes a

theyadoreshayyy:

@snowflake0531

snowflake0531:

It's not cosine

new2luv:

@snowflake0531 wrote:
It's not cosine
would it be c

theyadoreshayyy:

so is it c ?

snowflake0531:

\[\sin (45) = \frac {7}{AB} ~~~ \sin(45)AB = 7~~~AB= \frac{7}{\sin(45)}\]